OpenStudy (anonymous):

Nelson opened a savings account 11 years ago with a deposit of $3,115.89. The account has an interest rate of 5.3% compounded monthly. How much interest has Nelson earned?

OpenStudy (anonymous):

$2,458.77

OpenStudy (anonymous):

X = Initial Deposit P = Yearly Intrest Rate n = Number of Years Then Y = His Money after n Years = X(1+P/12)^12n = 3115.89(1+0.053/12)^12*11 = 5574.66 => The interest Nelson earned = Y - x = $2,458.77